# Triplet superconductivity by the orbital Rashba effect at surfaces of elemental superconductors

### v2

@@ Abstract
It is often assumed that in a superconductor without spin-triplet pairing, the formation of unconventional spin-triplet densities requires the spin-orbit interaction in combination with either broken inversion symmetry or broken time-reversal symmetry. Here, we show from first principles the existence of supercurrent-driven spin triplet densities on the surface of a variety of simple superconducting materials without spin-orbit coupling. We are able to attribute this phenomenon to the superconducting non-relativistic orbital Rashba Edelstein effect. Furthermore, we find that the spin-orbit induced spin moment is one order of magnitude smaller than the orbital moment, and has a vanishing effect on the total magnitude of the induced triplet density. Our findings imply the existence of a route to generate spin-currents without the use of heavy metals. Additionally, as an orbital moment can couple directly to a magnetic field, it shows that orbital physics is the dominant term that drives the superconducting diode effect.
@@
Recently, it has been suggested that there is an effect more fundamental to spin-Rashba splitting, which exists in inversion-broken systems irrespective of spin-orbit coupling. The orbital Rashba effect Park et al. (2011, 2012, 2013); Go et al. (2017, 2021a, 2021b) relies on the presence of multiple orbital degrees of freedom, a feature present in the majority of material surfaces. It has been shown to lead to a dramatic enhancement of the current-induced orbital magnetization as compared to its spin counterpart in experimental studies of the spin-orbit torques of Cu films Ding et al. (2020, 2022); Go et al. (2021a). The key to this enhancement is the so-called orbital Rashba-Edelstein effect mediated by strong orbital character of the states at the surface acquired as a result of broken inversion symmetry. Meanwhile, theoretical models have shown that this effect can be further enhanced when combined with superconductivity Chirolli et al. (2022); Ando et al. (2024). Experimental investigations of superconducting orbitronics are limited. However, experiments on superconducting spintronics have shown a dramatic enhancement of the signal when the metal undergoes the superconducting transition Linder and Robinson (2015); Jeon et al. (2020). This implies that superconducting orbitronics might lead to even further enhancement.
@@
$H_{\mathrm{DBdG}}(\mathbf{r})=\left(\begin{matrix}H_{D}(\mathbf{r})&\Delta_{\mathrm{eff}}(\mathbf{r})\\ \Delta^{*}_{\mathrm{eff}}(\mathbf{r})&-H^{*}_{D}(\mathbf{r})\end{matrix}\right),$ (1)
where $H_{D}(\mathbf{r})$ is the $4\times 4$ Dirac Hamiltonian and $\Delta_{\mathrm{eff}}(\mathbf{r})$ denotes the $4\times 4$ effective pairing potential of the superconductor which in this instance is considered spin singlet. Despite this, we calculate the three triplet components [$n^{\uparrow\uparrow}(\epsilon)$, $n^{\downarrow\downarrow}(\epsilon)$ and $n^{T_{z}}(\epsilon)=n^{\uparrow\downarrow}(\epsilon)+n^{\downarrow\uparrow}(\epsilon)$] of the density alongside the singlet component [$n^{s}(\epsilon)=n^{\uparrow\downarrow}(\epsilon)-n^{\downarrow\uparrow}(\epsilon)$] as they are not trivially zero when combining singlet superconductivity with magnetism, spin-orbit coupling and broken inversion symmetry Sato and Ando (2017). This first principles treatment of superconductivity has been used to model unconventional pairing Csire et al. (2018b); Ghosh et al. (2020), Rashba superconductivity Rüßmann and Blügel (2022a); Rüßmann et al. (2023), Iron-based superconductivity Reho et al. (2024a) and topological superconductivity Nyári et al. (2023); Lászlóffy et al. (2023) induced from impurities Saunderson et al. (2020b, 2022); Wu et al. (2023) and superconductor/topological insulator interfaces Rüßmann and Blügel (2022b); Reho et al. (2024b); Park et al. (2020). Further details of the first principles treatment of superconductivity can be found in the supplementary information Sup .
The setup is displayed in the schematic in Fig. 1. We model the effect of a supercurrent on a non-centrosymmetric superconductor by breaking the symmetry in the $x$-direction. The non-centrosymmetric superconductors chosen for this work are the surfaces of elemental superconductors. With the application of a current along $x$, the induced orbital polarization will be along $y$.
@@
We first compute the magnitude of the $n^{\uparrow\uparrow}(\epsilon)$ triplet density both without the spin-orbit coupling in Fig. 2(a) and with spin-orbit coupling on in Fig. 2(b). Remarkably, triplet densities are induced for each material in Fig. 2(a). When switching on spin-orbit coupling the most significantly enhanced triplet density in Fig. 2(b) is Pb, and the least is Al. This is consistent with the fact that SOC is much stronger in Pb than in Al, but rejects the conventional understanding that SOC is required to generate triplet density. To understand the microscopic origin of the spectrum shown in Fig. 2(a) we present the $y$ component of the orbital moment as a function of band filling $L_{y}(\epsilon)$ in Fig. 2(c). Here we see a remarkable correlation with the triplet density in Fig. 2(a) implying that the orbital moment plays the crucial role in determining the strength of the induced triplet density.
Figure 2: Current-induced imaginary part of $n^{\uparrow\uparrow}(\epsilon)$ triplet (a-b) and orbital (c) density around the superconducting gap $\Delta$ for Pb (blue), Al (red) and V (yellow). (a) The current-induced $n^{\uparrow\uparrow}(\epsilon)$ triplet density around the superconducting gap $\Delta$ for Pb (blue), Al (red) and V (yellow) with spin-orbit coupling. (b) Same as in (a) without spin-orbit coupling. (c) The $y$-component of the current-induced orbital ($L_{y}$) moment around the superconducting gap for Pb (blue), Al (red) and V (yellow) without spin-orbit coupling.
@@
To understand the significance of this result we discuss two experimental realizations which will benefit from and be explained by this phenomena.
@@
The second experiment is the superconducting diode effect Ando et al. (2020); Gutfreund et al. (2023); Nadeem et al. (2023). To perform this experiment, one sweeps a current in the ±$\pm x$ direction of the sample while keeping a constant magnetic field applied in the $y$ direction. When the material is a superconducting diode, a supercurrent will flow in the $+x$ direction, and a resistive current will flow in $-x$. The effect relies on the superconductor breaking inversion symmetry, giving rise to a global $y$ spin-polarization which breaks the symmetry of a ±$\pm x$ current sweep. Our findings show that the $y$ polarized orbital moment is significantly larger, even in Pb, than the spin moment. As an orbital moment can couple directly to the magnetic field it has its own diode effect which is distinct, and significantly more prominent, than the spin response. In the seminal work by F. Ando et al Ando et al. (2020), the superconducting diode effect was observed in Nb/Ta/V multilayers. We suggest an alternative heterostructure which still breaks inversion symmetry while minimizing the spin-orbit interaction such as Ti/Al/V, as we believe that this device would show the same diode effect just as strongly.
@@
To conclude, in this work we have shown that there is a coupling channel which exists between the orbital Rashba Edelstein effect and spin triplet Cooper pairs which circumnavigates spin Rashba spin-orbit coupling. This theoretical investigation was performed by modelling a supercurrent on the surface of the realistic superconductors V, Al and Pb. We show that while the spin-Rashba effect is present, the orbital Rashba effect is the dominant quantity which induces spin triplet densities even without spin-orbit coupling. Furthermore, the current-induced orbital moment is consistently higher than the current-induced spin moment. Finally, two experimental setups were discussed that could realise this theoretical discovery. Such discoveries open up a new chapter of unconventional superconductivity. One such example would be the question of orbital topology and its interaction with superconductivity would generate new forms of topological superconductivity which do not require spin-orbit coupling. Another such example would be the interaction with chiral phonons and crystals whose higher orbital angular momentum gives rise to superconductivity with higher degrees of freedom. Finally, this discovery provides new ways to incorporate superconductors into hybrid structures with magnetic systems, providing the blueprint for the next generation of computational architectures.
@@
Figure 4: A schematic illustration of nonlocal generation of SOTs in FM/SC/CuOx structures utilizing the superconducting OREE effect. The OREE effect from CuOx will generate triplet states in the superconductor, which will inject spin currents into the Ferromagnet via Andreev reflection.
@@ Acknowledgements.
The authors appreciate fruitful discussions with Prof. Balázs Ujfalussy, Dr. Philipp Rüßmann and Dr. Gabor Csire. This work was supported by the EIC Pathfinder OPEN grant 101129641 “OBELIX”. We also acknowledge the funding by the Deutsche Forschungsgemeinschaft TRR 173/3—268565370 (Projects A01 and A11), TRR 288—422213477 (Project B06). We gratefully acknowledge the Jülich Supercomputing Centre for providing computational resources under project jiff40. TGS gratefully acknowledges funding from TopDyn - Dynamics and Topology early career researchers grant, and acknowledges support from the US National Science Foundation (NSF) Grant Number 2201516 under the Accelnet program of Office of International Science and Engineering (OISE).

### v3

@@ Abstract
It is often assumed that in a superconductor without spin-triplet pairing, the formation of unconventional spin-triplet densities requires the spin-orbit interaction in combination with either broken inversion symmetry or broken time-reversal symmetry. Here, we show from first principles the existence of supercurrent-driven spin triplet densities on the surface of a variety of simple superconducting materials independent of spin-orbit coupling. We are able to attribute this phenomenon to the superconducting non-relativistic orbital Rashba Edelstein effect. Furthermore, we find that the spin-orbit induced spin moment is one order of magnitude smaller than the orbital moment, and has a vanishing effect on the total magnitude of the induced triplet density. Our findings imply the existence of a route to generate spin-currents without the use of heavy metals. Additionally, as an orbital moment can couple directly to a magnetic field, it shows that orbital physics is the dominant term that drives the superconducting diode effect.
@@
Recently, it has been suggested that there is an effect more fundamental to spin-Rashba splitting, which exists in inversion-broken systems irrespective of spin-orbit coupling. The orbital Rashba effect Park et al. (2011, 2012, 2013); Go et al. (2017, 2021a, 2021b) relies on the presence of multiple orbital degrees of freedom, a feature present in the majority of material surfaces. It has been shown to lead to a dramatic enhancement of the current-induced orbital magnetization as compared to its spin counterpart in experimental studies of the spin-orbit torques of Cu films Ding et al. (2020, 2022); Go et al. (2021a). The key to this enhancement is the so-called orbital Rashba-Edelstein effect mediated by strong orbital character of the states at the surface acquired as a result of broken inversion symmetry. Meanwhile, theoretical models have shown that this effect can be further enhanced when combined with superconductivity Chirolli et al. (2022); Ando et al. (2024). Experimental investigations of superconducting orbitronics are limited. However, experiments on superconducting spintronics have shown a dramatic enhancement of the signal when the metal undergoes the superconducting transition Linder and Robinson (2015); Jeon et al. (2020). This implies that superconducting orbitronics might lead to even further enhancement. Orbital moments and their connection to superconductivity have been investigated before Robbins et al. (2020) and it is clear that understanding this interaction will be fundamental for unconventional superconductivity.
@@
$H_{\mathrm{DBdG}}(\mathbf{r})=\left(\begin{matrix}H_{\mathrm{D}}(\mathbf{r})&\Delta_{\mathrm{eff}}(\mathbf{r})\\ \Delta^{*}_{\mathrm{eff}}(\mathbf{r})&-H^{*}_{\mathrm{D}}(\mathbf{r})\end{matrix}\right),$ (1)
where $H_{\mathrm{D}}(\mathbf{r})$ is the $4\times 4$ Dirac Hamiltonian and $\Delta_{\mathrm{eff}}(\mathbf{r})$ denotes the $4\times 4$ effective pairing potential of the superconductor which in this instance is considered spin singlet. Despite this, we calculate the three triplet components [$n^{\uparrow\uparrow}(\epsilon)$, $n^{\downarrow\downarrow}(\epsilon)$ and $n^{T_{z}}(\epsilon)=n^{\uparrow\downarrow}(\epsilon)+n^{\downarrow\uparrow}(\epsilon)$] of the density alongside the singlet component [$n^{s}(\epsilon)=n^{\uparrow\downarrow}(\epsilon)-n^{\downarrow\uparrow}(\epsilon)$] as they are not trivially zero when combining singlet superconductivity with magnetism, spin-orbit coupling and broken inversion symmetry Sato and Ando (2017). This first principles treatment of superconductivity has been used to model unconventional pairing Csire et al. (2018b, 2022); Ghosh et al. (2020), Rashba superconductivity Rüßmann and Blügel (2022a); Rüßmann et al. (2023), Josephson Junctions Yamazaki et al. (2024), Iron-based superconductivity Reho et al. (2024a) and topological superconductivity Nyári et al. (2023); Lászlóffy et al. (2023) induced from impurities Saunderson et al. (2020b, 2022); Wu et al. (2023) and superconductor/topological insulator interfaces Rüßmann and Blügel (2022b); Reho et al. (2024b); Park et al. (2020). Further details of the first principles treatment of superconductivity can be found in the supplementary information Sup .
The induced supercurrent can be modelled via a modulation of the phase of the superconducting order paramter, such as in a Josephson Junction Yamazaki et al. (2024). In this instance the symmetry breaking effect of a supercurrent can be modelled as a linear phase gradient Chirolli et al. (2022), where the setup is displayed in the schematic in Fig. 1. We model the effect of a supercurrent on a non-centrosymmetric superconductor by modulating the phase of the order parameter along the $x$-direction. The non-centrosymmetric superconductors chosen for this work are the surfaces of elemental superconductors. With the application of a current along $x$, the induced orbital polarization will be along $y$.
@@
We first compute the magnitude of the $n^{\uparrow\uparrow}(\epsilon)$ triplet densities for the selected materials in Fig. 2(a). As expected, triplet densities are induced for each material in Fig. 2(a) which remarkably seem to be irrespective of spin-orbit strength. Scaling the spin-orbit parameter to 1% in Fig. 2(b) we see that the triplet density is most greatly affected in Pb, and the least affected in Al. This is consistent with the fact that SOC is much stronger in Pb than in Al, but rejects the conventional understanding that SOC will tune the strength of the effect. To understand the microscopic origin of the spectrum shown in Fig. 2(b) we present the $y$ component of the orbital $L_{y}(\epsilon)$ and spin $S_{y}(\epsilon)$ moments as a function of band filling in Fig. 2(c) and (d). Firstly, the scale of figure (d) is two orders of magnitude smaller than (c), which is the expected result with such a small spin orbit strength, and implies that its influence on the triplet density is immaterial. Comparing Fig. 2(b) with (c) we see a strong correlation implying that the orbital moment plays the crucial role in determining the strength of the induced triplet density in the limit of low spin orbit strength. This almost implies that there are effectively two channels to that drive the triplet density in noncentrosymmetric superconductors.
Figure 2: Current-induced imaginary part of $n^{\uparrow\uparrow}(\epsilon)$ triplet (a-b) and orbital (c) density around the superconducting gap $\Delta$ for Pb (blue), Al (red) and V (yellow). (a) The current-induced $n^{\uparrow\uparrow}(\epsilon)$ triplet density around the superconducting gap $\Delta$ for Pb (blue), Al (red) and V (yellow) with spin-orbit coupling. (b) Same as in (a) with only 1% spin-orbit coupling strength. (c) The $y$-component of the current-induced orbital ($L_{y}$) moment around the superconducting gap for Pb (blue), Al (red) and V (yellow) with 1% spin-orbit coupling. (d) The $y$-component of the current-induced spin ($S_{y}$) moment around the superconducting gap for Pb (blue), Al (red) and V (yellow) with 1% spin-orbit coupling.
@@
The second experiment is the superconducting diode effect Ando et al. (2020); Gutfreund et al. (2023); Nadeem et al. (2023). To perform this experiment, one sweeps a current in the ±$\pm x$ direction of the sample while keeping a constant magnetic field applied in the $y$ direction. The Rashba textures present in the system induce a y-polarized moment as exemplified in this work. The constant applied magnetic field therefore acts as a symmetry breaking force which interacts with the induced moment which will change sign depending on whether the current is swept in the $+x$ or the $-x$ direction. The result will be nonreciprocal behaviour between the critical current in the $+x$ direction versus the critical current in the $-x$ direction. This effect relies on the superconductor breaking inversion symmetry, giving rise to a global $y$ spin-polarization which breaks the symmetry of a ±$\pm x$ current sweep. Our findings show that the $y$ polarized orbital moment is significantly larger, even in Pb, than the spin moment. As an orbital moment can couple directly to the magnetic field it has its own diode effect which is distinct, and significantly more prominent, than the spin response. This means that, whilst the current experimental search has been to increase the atomic number in order to enhance the spin-orbit coupling, we suggest that this is not the most important effect. For example in the seminal work by F. Ando et al Ando et al. (2020), the superconducting diode effect was observed in Nb/Ta/V multilayers, whilst we suggest that in fact light elemental superconductors will already be sufficient.
@@
To conclude, in this work we have shown that there is a coupling channel which exists between the orbital Rashba Edelstein effect and spin triplet Cooper pairs which circumnavigates spin Rashba spin-orbit coupling. This theoretical investigation was performed by modelling a supercurrent on the surface of the realistic superconductors V, Al and Pb. We show that while the spin-Rashba effect is present, the orbital Rashba effect is the dominant quantity which induces spin triplet densities even 1% spin-orbit coupling. Furthermore, the current-induced orbital moment is consistently higher than the current-induced spin moment. Finally, two experimental setups were discussed that could realise this theoretical discovery. Such discoveries open up a new chapter of unconventional superconductivity. One such example would be the question of orbital topology and its interaction with superconductivity would generate new forms of topological superconductivity. Another such example would be the interaction with chiral phonons and crystals whose higher orbital angular momentum gives rise t o superconductivity with higher degrees of freedom. Finally, this discovery provides new ways to incorporate superconductors into hybrid structures with magnetic systems, providing the blueprint for the next generation of computational architectures.
@@
Figure 4: A schematic illustration of nonlocal generation of SOTs in FM/SC structures utilizing the superconducting OREE effect at the interface between the SC and FM. The OREE effect from the interface will generate triplet states in the superconductor, which will in turn be injected into the Ferromagnet.
@@ Acknowledgements.
The authors appreciate fruitful discussions with Prof. Balázs Ujfalussy, Dr. Philipp Rüßmann and Dr. Gabor Csire. This work was supported by the EIC Pathfinder OPEN grant 101129641 “OBELIX”. We also acknowledge the funding by the Deutsche Forschungsgemeinschaft TRR 173/3—268565370 (Projects A01 and A11), TRR 288—422213477 (Project B06). We gratefully acknowledge the Jülich Supercomputing Centre for providing computational resources under project jiff40. TGS gratefully acknowledges funding from TopDyn - Dynamics and Topology early career researchers grant, and acknowledges support from the US National Science Foundation (NSF) Grant Number 2201516 under the Accelnet program of Office of International Science and Engineering (OISE). This material is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-23-1-0132.